# Uncertainty Matters in Dynamic Gaussian Splatting for Monocular 4D Reconstruction

### v1

@@ 6 Discussion and Conclusion
We introduce a novel dynamic Gaussian Splatting framework USplat4D, showing that uncertainty matters in dynamic Gaussian splatting. We propose time-varying uncertainty estimation method and built an uncertainty-guided graph, systematically demonstrating improved geometric consistency and rendering quality under extreme views. While our method still inherits errors from the underlying Gaussian splatting model, we hope that our findings could encourage further work on leveraging uncertainty to advance robust 4D reconstruction. Please see appendix for more discussion.
@@
This appendix provides additional theoretical derivations, implementation details, experimental results, ablation studies, and discussions on broader impacts. A quick outline is as follows:
  • Section A: Additional details on our proposed USplat4D.
    • Section A.1: Derivation of the uncertainty model.
    • Section A.2: Details of the training losses.
  • Section B: Implementation details, including adaptation of base models and training schedules.
    • Section B.1: Adaptation of SoM and MoSca.
    • Section B.2: Training USplat4D Model.
  • Section C: Additional experimental results.
    • Section C.1: Comparison with additional NeRF-based methods on the DyCheck dataset.
    • Section C.2: Tracking results on the DyCheck dataset.
    • Section C.3: Per-scene results on the DyCheck dataset.
    • Section C.4: Data curation for the Objaverse dataset.
    • Section C.5: Per-scene results on the Objaverse dataset.
    • Section C.6: Results on the NVIDIA dataset.
  • Section D: Additional ablation studies and analyses.
    • Section D.1: Ablation study on key / non-key ratio.
    • Section D.2: Analysis on key node selection.
    • Section D.3: Analysis on runtime.
    • Section D.4: Failure cases.
  • Section E: Discussion of the social and broader impact of our work.
@@ A.2.1 Motion Loss
The vanilla Gaussian splatting approach (e.g., Luiten et al. (2024)) represents a dynamic scene using 3D Gaussians with time-varying motion and define motion loss. The recent paper also follow up on the motion loss and modifies it based on their tasks. Here, we give the general form of these motion losses. The isometry loss describing the distance constraint between Gaussians is defined as
$\mathcal{L}^{\text{iso}}=\frac{1}{k|\mathcal{V}|}\sum_{t=0}^{T-1}\sum_{i\in\mathcal{V}}\sum_{j\in\text{knn}_{i;k}}w_{i,j}(\|\mathbf{p}_{j,n}-\mathbf{p}_{i,n}\|_{2}-\|\mathbf{p}_{j,t}-\mathbf{p}_{i,t}\|_{2})$ (S8)
where subscript $n$ means canonical space, $\mathbf{t}$ is the Gaussian’s 3D position vector, and $w_{i,j}$ represent the edge weight between $i$-th and $j$-th Gaussians. Since isometry does not take coordinates into consideration, we use rigidity loss to unify the coordinates and constrain the relative motion between Gaussians. The rigidity loss is defined by
$\mathcal{L}^{\text{rigid},\Delta}=\frac{1}{k|\mathcal{V}|}\sum_{t=\Delta}^{T-1}\sum_{i\in\mathcal{V}}\sum_{j\in\text{knn}_{i;k}}w_{i,j}\left\|\left(\mathbf{p}_{j,t-\Delta}-\mathbf{T}_{i,t-\Delta}\mathbf{T}_{i,t}^{-1}\mathbf{p}_{j,t}\right)\right\|_{2}$ (S9)
@@ B.2 Training USplat4D Model
During preprocessing of USplat4D, we sequentially initialize the key and non-key graphs. To train the USplat4D model, we build on pretrained base models and allocate additional training iterations for fair comparison. Specifically, we train SoM for 400 extra epochs and MoSca for 1600 extra steps, using a consistent batch size of 8. These schedules are empirically chosen to ensure good convergence while aligning with the respective optimization routines of the base models.
For the first 10% and the last 20% of the training duration, we disable both density control and opacity reset to maintain stability. For the remaining iterations, we enable density control and opacity reset to improve rendering quality. The per-epoch and per-step training time is similar to that of SoM and MoSca, respectively.
@@ Appendix C Additional Experimental Results
5 scenes
1 $\times$ resolution
DynIBaR (Li et al., 2023) 13.41 0.48 0.55
@@ 1 $\times$ resolution
7 scenes
2 $\times$ resolution
D-NPC (Kappel et al., ) 16.41 0.582 0.319
@@ C.1 Comparison with Additional NeRF-based Methods on DyCheck Dataset
In Table 1 of the main paper, we primarily compare with Gaussian Splatting based methods that use explicit 3D representations. Here, we extend the comparison by including additional NeRF-based methods (Miao et al., 2024; Park et al., 2021; Gao et al., 2022; Li et al., 2023; Sun et al., 2024a; Kappel et al., ; Zhao et al., 2024) that adopt implicit neural radiance fields. While these methods are effective for static or mildly dynamic scenes, our experiments show that Gaussian splatting based approaches consistently outperform NeRF-based methods on the DyCheck dataset (Gao et al., 2022). This comparison serves as an extension of Table 1 and the results are summarized in Table S1.
@@ C.6 Results on NVIDIA Dataset
Our main quantitative analysis focuses on the DyCheck-iPhone dataset (Gao et al., 2022) in the main paper, which aligns well with our goal of synthesizing dynamic objects captured by a moving camera (see Section 1 of the main paper). Here we also evaluate our method on the NVIDIA dataset (Yoon et al., 2020) for broader comparison by following the evaluation protocal in RDRF (Liu et al., 2023). The results are shown in Table S5.
We note that USplat4D also improves performance when applied to state-of-the-art MoSca (Lei et al., 2025) on the NVIDIA dataset, although the gains are marginal. This is expected, as the NVIDIA dataset (or other datasets such as HyperNeRF (Park et al., 2021) dataset) differs from DyCheck to feature input views with more limited motion.
Table S5: Results on the NVIDIA dataset Yoon et al. (2020).
@@ Appendix D Additional Ablation and Analysis
D.1 Ablation Study on Key / Non-Key Ratio
@@
D.2 Analysis on Key Node Selection
In Section 4.2, we design the key node selection strategy by sampling key node candidates via 3D gridization to ensures spatial coverage and reduces redundancy. To test its functionality, we replace it by a random sampling strategy. Specifically, we keep the number of selected Gaussians per frame the same, but sample them randomly from all remaining grids without enforcing per-grid constraints. Table S7 demonstrates the performance drops on the DyCheck validation set, which causes by a non-uniform spatial distribution (i.e., some grids may contain multiple selected nodes, while some are missed entirely).
Table S7: Ablation study on key node selection strategy. We assess the impact of the key node sampling strategy by randomly sampling without considering spatial distribution.
@@ D.2 Analysis on Key Node Selection
USplat4D (full model) 19.63 0.716 0.249
   w/o spatial-random key node selection 19.50 0.712 0.251
D.3 Analysis on Runtime
@@
D.4 Limitations and Failure Cases
Although the model partially relies on the tracking results of the initial 4D Gaussians and the geometry information during the observation time period, both quantitative and qualitative results demonstrate the robustness of the model in restoring or preserving shape when faced with partial observation and tracking failure. However, the graph model still suffers from the incorrect motion in the 4D prior. Although our model can utilize uncertainty model to distinguish the usefully information but if the 4D prior totally fails without leaving useful information, our model can not retrieve useful information from it. According to our observations, when the video only includes few frames with large motion and the initial 4D Gaussians do not provide stable tracking results, the output of the USplat4D may not guarantee the geometry consistency. The failure case can be found in the demo video.
Appendix E Social and Broader Impact
@@
Our work advances dynamic 3D scene reconstruction and novel view synthesis from monocular videos, with applications in AR/VR, physical scene understanding, digital content creation, and human-object interaction. By modeling uncertainty and improving synthesis under extreme viewpoints, our method contributes to more robust and accessible 4D modeling. While the method has positive use cases, it could be misused for synthetic content manipulation. Our method does not involve sensitive data, and care should be taken in downstream applications to ensure responsible use.

### v2

@@ 6 Discussion and Conclusion
We introduce USplat4D, a novel dynamic Gaussian Splatting framework that demonstrates that how uncertainty estimation is pivotal for monocular 4D reconstruction. By integrating time-varying uncertainty estimation with an uncertainty-guided graph, our approach significantly improves geometric consistency and rendering quality, particularly under challenging extreme novel view synthesis. While our framework remains influenced by the computational overhead and inherent errors of underlying visual foundation models, we hope that our findings could encourage further work on leveraging uncertainty to advance robust 4D reconstruction and extreme novel view synthesis.
@@ A.2.1 Motion Loss
The vanilla Gaussian splatting approach (e.g., (Luiten et al., 2024)) represents a dynamic scene using 3D Gaussians with time-varying motion and define motion loss. The recent paper also follow up on the motion loss and modifies it based on their tasks. Here, we give the general form of these motion losses. The isometry loss describing the distance constraint between Gaussians is defined as
$\mathcal{L}^{\text{iso}}=\frac{1}{k|\mathcal{V}|}\sum_{t=0}^{T-1}\sum_{i\in\mathcal{V}}\sum_{j\in\text{knn}_{i;k}}w_{i,j}(\|\mathbf{p}_{j,o}-\mathbf{p}_{i,o}\|_{2}-\|\mathbf{p}_{j,t}-\mathbf{p}_{i,t}\|_{2})$ (S8)
where subscript $o$ means canonical space, $\mathbf{t}$ is the Gaussian’s 3D position vector, and $w_{i,j}$ represent the edge weight between $i$-th and $j$-th Gaussians. Since isometry does not take coordinates into consideration, we use rigidity loss to unify the coordinates and constrain the relative motion between Gaussians. The rigidity loss is defined by
$\mathcal{L}^{\text{rigid},\Delta}=\frac{1}{k|\mathcal{V}|}\sum_{t=\Delta}^{\mathbf{T}-1}\sum_{i\in\mathcal{V}}\sum_{j\in\text{knn}_{i;k}}w_{i,j}\left\|\left(\mathbf{p}_{j,t-\Delta}-\mathbf{T}_{i,t-\Delta}\mathbf{T}_{i,t}^{-1}\mathbf{p}_{j,t}\right)\right\|_{2}$ (S9)
@@ B.2 Training USplat4D Model
During preprocessing of USplat4D, we sequentially initialize the key and non-key graphs. To train the USplat4D model, we build on pretrained base models and allocate additional training iterations for fair comparison. Specifically, we train SoM for 400 extra epochs and MoSca for 1600 extra steps, using a consistent batch size of 8. These schedules are empirically chosen to ensure good convergence while aligning with the respective optimization routines of the base models. For the first 10% and the last 20% of the training duration, we disable both density control and opacity reset to maintain stability. For the remaining iterations, we enable density control and opacity reset to improve rendering quality. The per-epoch and per-step training time is similar to that of SoM and MoSca, respectively.
@@ Appendix C Additional Experimental Results
5 scenes 1 $\times$ resolution DynIBaR (Li et al., 2023) 13.41 0.48 0.55
@@
7 scenes 2 $\times$ resolution D-NPC (Kappel et al., 2025) 16.41 0.582 0.319
@@ C.1 Comparison with Additional NeRF-based Methods on DyCheck Dataset
In Table 1 of the main paper, we primarily compare with Gaussian splatting based methods that use explicit 3D representations. Here, we extend the comparison by including additional NeRF-based methods (Miao et al., 2024; Park et al., 2021; Gao et al., 2022; Li et al., 2023; Sun et al., 2024a; Kappel et al., 2025; Zhao et al., 2024) that adopt implicit neural radiance fields. While these methods are effective for static or mildly dynamic scenes, our experiments show that Gaussian splatting based approaches consistently outperform NeRF-based methods on the DyCheck dataset (Gao et al., 2022). This comparison serves as an extension of Table 1 and the results are summarized in Table S1.
@@ C.6 Results on NVIDIA Dataset
Our main quantitative analysis focuses on the DyCheck-iPhone dataset (Gao et al., 2022) in the main paper, which aligns well with our goal of synthesizing dynamic objects captured by a moving camera (see Section 1 of the main paper). Here we also evaluate our method on the NVIDIA dataset (Yoon et al., 2020) for broader comparison. We follow the evaluation pipeline in RoDynRF (Liu et al., 2023). The results are shown in Table S5. We note that USplat4D also improves performance when applied to state-of-the-art MoSca (Lei et al., 2025) on the NVIDIA dataset, although the gains are marginal. This is expected, as the NVIDIA dataset (or other datasets such as HyperNeRF (Park et al., 2021) dataset) differs from DyCheck to feature input views with more limited motion. In contrast, datasets with larger camera movement or more challenging dynamics (DyCheck and Objaverse) benefit more from our design, leading to more substantial improvements.
Table S5: Results on the NVIDIA dataset (Yoon et al., 2020).
@@
Figure S2: Comparison on HyperNeRF dataset. The groundtruth, 4DGS, and USplat4D (ours) are arranged from left to right. The red and green boxes highlights regions with fine-grained details. The values displayed in the top-right corner represent PSNR, SSIM, and LPIPS, arranged from top to bottom. Our method reconstructs high-frequency structures more faithfully and preserves finer geometry compared with both 4DGS.
C.7 Comparison on HyperNeRF dataset
We further provide a qualitative comparison among 4DGS (Wu et al., 2024), and USplat4D (ours) on the HyperNeRF (Park et al., 2021) dataset, as shown in Figure S2. By focusing on the dynamic, detail-rice regions, we observe that USplat4D restores noticeably finer details and sharper high-frequency structures. While 4DGS tends to oversmooth or lose small geometric components, our uncertainty-guided graph model consistently preserves these structures, leading to more stable and detailed 4D reconstructions.
@@ Appendix D Additional Ablation and Analysis
D.1 Ablation study on key / non-key ratio.
@@
D.2 Ablation study on color threshold
Table S7: Ablation study on the color threshold $\eta_{c}$ using Objaverse dataset. Results are reported relative to our chosen threshold ($\eta_{c}=0.50$). $\eta_{c}=1.00$ means no thresholding.
@@ D.2 Ablation study on color threshold
MoSca (baseline) 0.980 0.992 1.115
0.01 0.993 0.999 1.023
0.10 0.996 1.000 1.011
0.20 0.994 1.000 1.014
0.40 0.998 1.000 1.008
0.50 (reference) 1.000 1.000 1.000
0.80 1.001 1.000 1.001
1.00 0.997 1.000 1.010
The color threshold $\eta_{c}$ (see Equation 6) is introduced to prevent incorrect uncertainty estimation from Equation 3 when the prior Gaussians have not yet converged on certain pixels in the input images. We further study the influence of $\eta_{c}$ on the Objaverse dataset. As reported in Table S7, performance degrades when $\eta_{c}$ is set to $1.0$, which corresponds to disabling the threshold entirely. In this case, the model incorrectly trusts Gaussians with large color errors and assigns them falsely low uncertainty. On the other hand, when $\eta_{c}$ is reduced below $0.1$, many genuinely low-uncertainty Gaussians are mistakenly treated as unreliable, which again harms reconstruction quality. Although these two extremes lead to worse performance, we find a broad performance plateau between $[0.4,0.8]$, showing that the hyperparameter is easy to select and not sensitive in practice. Across all tested values of $\eta_{c}$, our approach consistently outperforms MoSca.
D.3 Ablation Study on the Significant Period Threshold
As discussed in Section 4.2 of the main paper, we use a hyper-parameter (namely, Significant Period Threshold or SPT) to filter out Gaussians that are well observed for only a short duration (no more than four frames). Such transient Gaussians contribute little to motion propagation and often lead to weak or unreliable graph connections. In our experiments, we uniformly set the SPT to 5 for all datasets. To further assess its influence, we evaluate several values of SPT, as shown in Table S8. We observe a broad performance plateau when $\textrm{SPT}\geq 3$, indicating that the model is stable across a reasonable range of choices. Performance drops when the threshold is set to $1$ (i.e., disabling the filtering by SPT) which introduces short-lived Gaussians into the key graph, which matches the intuition described above. For all choices of SPT, our method consistently achieves higher performance than MoSca across all tested values.
Table S8: Ablation study on the Significant Period (SPT). Results are reported relative to our chosen significant period threshold ($\text{SPT}=5$).
@@ D.3 Ablation Study on the Significant Period Threshold
MoSca (baseline) 0.980 0.992 1.109
1 (i.e., no thresholding) 0.994 0.999 1.016
3 0.998 1.000 1.006
5 1.000 1.000 1.000
7 0.996 1.000 1.007
10 0.996 1.000 1.004
D.4 Ablation Study on the Scaling Factor
The scaling factor $r_{x},r_{y},r_{z}]$ controls the relative weight when transforming 2D uncertainty into axis-aligned 3D uncertainty. The key insight is that the reliability of depth varies with camera motion. In scenarios where the camera undergoes large translation in the $x$-$y$ plane, the depth becomes better constrained. In such cases, down-weighting the depth component (using a smaller $r_{z}$) reduces noise in the uncertainty estimate. This intuition aligns with geometric models such as Direct Sparse Odometry (Engel et al., 2017), which models depth estimates with a variance term that is inversely proportional to the baseline, i.e., $\Delta Z\propto\frac{1}{b}$, where $\Delta Z$ is depth standard deviation and $b$ is camera baseline or translation.
Our design principle is that the default setting $[1,1,1]$ is valid, and once the camera motion is unbalanced across directions, adjusting the weights can improve robustness. For all experiments, we set $r_{x},r_{y},r_{z}]=[1,1,0.01]$. We also evaluated other values of $r_{z}$ and did not observe notable differences (a small value works consistently). For datasets such as DyCheck, DAVIS, and NVIDIA, which have limited camera movement along the depth axis and often emphasize rotation or small shifts, using a smaller $r_{z}$ improves the stability of the uncertainty estimation. For objaverse, where the camera moves across all three spatial directions, adjusting $r_{z}$ from $1.0$ to $0.01$ has minimal side effect, as shown Table S9.
Table S9: Ablation study on the depth uncertainty scale ratio $r_{z}$. Results are reported relative to our chosen ratio threshold (1.00).
[…]
MoSca (baseline) 0.989 0.992 1.100
0.01 1.001 0.999 1.005
1.00 (ours, reference) 1.001 1.000 1.000
@@ D.4 Ablation Study on the Scaling Factor
D.5 Analysis on Runtime
[…]
To better understand how the time usage increases with video length, we further curate training scenes with different sequence lengths from the Objaverse dataset. Figure S3 shows the runtime of MoSca and our method relative to the input video length. The runtime of our method exhibits a smaller slope in both stages compared with MoSca.
Figure S3: Time usage. MoSca involves a preprocessing stage followed by a reconstruction stage. MoSca Total is the sum of MoSca Precompute and MoSca Reconstruct. With the Gaussian priors from MoSca, our method includes a graph initialization stage (which covers uncertainty estimation and graph construction) and an overall training stage. USplat4D Total is the sum of Graph Initialization and Overll Training.
Figure S4: Failure of tracking on textureless surface. The input camera moves in a circular trajectory around a chick that remains quasi-static. We show the tracking and reconstruction results at frame 32 for both SoM and USplat4D (ours). Because the chick’s surface lacks texture, the tracks sampled at the initial frame (t = 0) drift and accumulate incorrect motion under SoM, eventually causing the reconstruction to collapse. In contrast, our method is able to partially recover the geometry and produce more stable tracking.
Figure S5: Comparison of reconstruction results on fast-motion objects. For each scene, the top row shows the overall view rendered from input views and novel views, and the three rows below highlight the fast-motion regions across ten successive time frames.
Figure S6: Qualitative on deforming objects. First row (Davis dancing scene): The subject’s arms exhibit strong non-rigid motion throughout the dance. Second row (DyCheck handwavy scene): The hand undergoes continuous pose changes with significant articulation. Third row (DyCheck mochi-high-five scene): The cat’s head rotates substantially. MoSca often produces drifting geometry where the ears or back of the neck separate from the head, while our method preserves consistent structure across time and viewpoints.
Appendix E Analysis of Challenging Cases
Although our model partially relies on the tracking quality of the initial 4D Gaussians and the geometric cues available during their observation period, both the quantitative and qualitative results demonstrate strong robustness in restoring or preserving shape under partial visibility and imperfect tracking. The uncertainty-guided graph allows the model to downweight unreliable motion and propagate stable cues across space and time. However, the framework still inherits the fundamental limitations of the underlying 4D prior. When the prior does not contain any reliable motion or geometric information, the uncertainty model has no meaningful signal to propagate, and therefore cannot fully recover the missing structures. Here, we further examine the impact of most common challenging cases: textureless regions, fast motion, and deformable objects.
E.1 Textureless Region
In textureless areas, the vision foundation model produces unreliable tracks, which in turn causes SoM and MoSca to generate incorrect dynamic Gaussian priors. When our model receives these flawed priors, the uncertainty-guided graph attempts to propagate geometric information under the guidance of uncertainty across time and space. However, when the underlying correct information is rare, the propagated motion inevitably suffers from the poor initial tracking and reconstruction. In the Objaverse experiments, such textureless regions noticeably reduce quantitative performance compared to real-world datasets with richer appearance cues. Under our circling-camera setup, a typical failure case is shown in Figure S4. The Gaussians from SoM drift with the camera instead of representing stable geometric structures, while our method stabilizes the overall motion but still inherits errors in local regions where tracking is fundamentally unreliable.
E.2 Fast Motion
Monocular 4D reconstruction under fast motion remains highly challenging, yet our method demonstrates strong robustness in such scenarios. By employing an uncertainty-weighted DQB loss together with trainable motion parameters $\mathbf{p}_{i,t}$, the model can deviate from node-based interpolation when uncertainty signals unreliable priors, enabling recovery of local fast motion even under incorrect initialization (Figure S5). This capability arises from relaxing the hard constraints imposed by node-based interpolation. In contrast, existing monocular 4D Gaussian splatting approaches, such as SoM and MoSca, rely heavily on motion regularizers that favor smooth, temporally coherent trajectories, implicitly biasing them toward slow or moderate motion. Their motion models further limit fast dynamics: SoM restricts motion to a small set of bases, capping allowable local deformation, while MoSca interpolates trajectories via DQB on scaffolds, producing stable but nearly rigid motion. When scaffolds are missing in high-motion regions, this interpolation becomes a hard constraint that prevents Gaussians from following rapid, input-aligned motion. Despite these improvements, our reconstructions in fast-motion regions still inherit rolling-shutter distortions from the input video and rapid motion further challenges the preservation of multi-view consistency. Addressing such artifacts remains an open problem in monocular 4D reconstruction and is beyond the scope of this work.
E.3 Deforming Objects
Considering that deformable objects pose greater challenges than near-rigid ones, we conduct a qualitative evaluation of monocular 4D reconstruction in highly deforming scenarios. Figure S6 compares novel-view synthesis results of our method and MoSca across several subjects. Our reconstructions align more faithfully with the underlying geometry, particularly in regions with large non-rigid motion. This robustness stems from our uncertainty-guided spatio-temporal graph and uncertainty-weighted DQB loss, which softly regularize non-key nodes while allowing deviations when priors are unreliable. Consequently, even with imperfect Gaussian initialization from MoSca, our framework recovers fine-scale deforming geometry while maintaining spatio-temporal consistency. In contrast, MoSca interpolates all Gaussians using scaffold-based skinning weights. While effective with sufficient scaffold coverage, this design becomes brittle when scaffolds fail to capture corners, extremities, or thin structures due to hyperparameter sensitivity or inaccurate 2D priors. Gaussians in these regions become overly constrained, leading to distortions or missing parts, as seen in highly deformable structures such as human limbs or the cat’s ear in Figure S6. Despite these improvements, reconstructing highly deformable regions remains challenging when priors are severely inaccurate or observations are sparse, indicating directions for future work.
Appendix F Social and Broader Impact
@@
Our work advances dynamic 3D scene reconstruction and novel view synthesis from monocular videos, with potential applications in AR/VR, physical scene understanding, digital content creation, and human-computer interaction. By modeling uncertainty and improving synthesis under extreme viewpoints, our method contributes to more robust and accessible 4D modeling. While the method has positive use cases, it could be misused for synthetic content manipulation. Our method does not involve sensitive data, and care should be taken in downstream applications to ensure responsible use.